When 25.0 mL of a 8.17×10-4 M zinc iodide solution is combined with 22.0 mL of a 7.60×10-4 M potassium hydroxide solution does a precipitate form?  (yes or no)

For these conditions the Reaction Quotient, Q, is equal to .
Answer #1

2949 The solubility product of Zn (OH)2 = 5x10-17 The final volume = (25+22) me = 47m2 So the concentration of Zn2+ ion in th